OpenStudy (lxelle):

How does ln(100-x)^e= (-0.02t + c)^e become 100-x= ce^0.02t?? I got the LHS, but not the RHS

OpenStudy (lxelle):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[\ln(100-x)^e= (-0.02t + c)^e\\ e\ln(100-x)= (-0.02t + c)^e\\ \ln(100-x)= (-0.02t + c)^e/e\\ 100-x= e^{(-0.02t + c)^e/e}\] ... hmm

OpenStudy (lxelle):

I don't get your rhs D:

OpenStudy (unklerhaukus):

maybe i made a mistake, what did you get?

OpenStudy (lxelle):

In a chemical reaction a compound X is formed from a compound Y. The masses in grams of X and Y present at time t seconds after the start of the reaction are x and y respectively. The sum of the two masses is equal to 100 grams throughout the reaction. At any time, the rate of formation of X is proportional to the mass of Y at that time. When t = 0, x = 5 and dx/dt = 1.9. (i) Show that x satisfies the differential equation: dx/dt = 0.02(100 − x). (ii) Solve this differential equation, obtaining an expression for x in terms of t. [6]

OpenStudy (lxelle):

Was trying to solve ii

OpenStudy (lxelle):

Fromm previous ques, Dx/dt =0.02 (100-x) I got ln (100-x)^e = (-0.02t + c)^e

OpenStudy (unklerhaukus):

\[dx/dt = 0.02(100-x)\] this the variables are separable \[dx/(100-x) = 0.02dt\] now integrate \[\int dx/(100-x) = \int 0.02\,dt\]

OpenStudy (lxelle):

Lol I alrd said I got ln(100-x)^e = (-0.02t +c)^e) then what's next

OpenStudy (michele_laino):

your equation is: \[\dot x + kx = 100k\] where k is such that: \[\dot x = ky\]

OpenStudy (lxelle):

@Michele_Laino are we even talking abt the same ques

OpenStudy (michele_laino):

yes! it is your question

OpenStudy (lxelle):

I'm talking abt ii). -.-

OpenStudy (michele_laino):

ok! I substitute k = 0.02 and I will integrate that equation

OpenStudy (lxelle):

How do you freaking integrate it lol. That's my main prob ugh

OpenStudy (unklerhaukus):

\[\int dx/(100-x) = \int 0.02\,dt\\ -\ln|100-x|=0.02t+c\]

OpenStudy (lxelle):

Uh huh

OpenStudy (michele_laino):

if you have to integrate this equation: \[\dot x + P(t)x = Q(t)\] then your solution x(t) is: \[\begin{gathered} x\left( t \right) = \exp \left( { - \int {P(t)dt} } \right)\left( {c + \int {Q(t)\exp \left( {\int {P(t)dt} } \right)dt} } \right) \hfill \\ x(0) = c \hfill \\ \end{gathered} \]

OpenStudy (lxelle):

@Michele_Laino Sorry but not sorry. You're totally not helping. I suggest you go help others. Thanks

OpenStudy (michele_laino):

I'm very sorry, but I suggest you that you have to study about how to integrate some first order ODE's @LXelle

OpenStudy (unklerhaukus):

\[\ln|100-x|=-0.02t+c\\ 100-x=e^{-0.02t-c}\\ 100-x=e^{-0.02t}/e^c\\ 100-x=Ce^{-0.02t}\]

OpenStudy (unklerhaukus):

first order technique is to general here, it is a variables separable DE

OpenStudy (michele_laino):

yes! you are right! @UnkleRhaukus nevertheless there are many ways in order to integrate a first order ODE (ordinary differential equation)

OpenStudy (unklerhaukus):

keep in mind, this is \(\textit{chemistry}\)

OpenStudy (michele_laino):

keep in mind that mathematics doesn't distinguish among physics, chemistry etc.. @UnkleRhaukus

OpenStudy (lxelle):

@Michele_Laino i expected you to explain and not just throw me some formulas. You weren't even trying to understand my problem

OpenStudy (michele_laino):

you are right @LXelle nevertheless my formula is a statement of a theorem

OpenStudy (lxelle):

Anyways thanks @UnkleRhaukus

OpenStudy (michele_laino):

please be more polite @LXelle thanks!

OpenStudy (lxelle):

Whatever

OpenStudy (unklerhaukus):

@LXelle, you could be a little more polite to @Michele_Laino,'s posts Are at least directly relevant to you question.